Suppose you have two RAID arrays, one implementing RAID 3, the other RAID 5. Each has 9 disk drives in its array. If the RAID 5 array can read a 1 Kbyte of data in 16 ms, how long would you expect the RAID 3 array to take to read 1 Kbyte of data? Why?
Answer #1

In the system we have two types of RAID (RAID 3 and RAID 5) with 9 disks. At high level following are the key charactestics of these RAID implementations:

RAID 3:

  • Stores data in different HDD.
  • Parity is stored in one HDD so all data srips in n-1 HDD and one HDD for Parity.
  • Byte wise striping
  • Party is bitwise.
  • Write Performance is poor and read is fair.

RAID 5:

  • Stores data in blocks in different HDD and all HDD are used for data.
  • Parity is distributed across all HDD.
  • Block wise stripping
  • Parity is Block wise.
  • Read data is very fast however write is slower as needs to calculate parity.

since in RAID 3 data is stripped bytewise and RAID 5 its blockwise (Assume 128 Bytes of block). Assuming all disk controllers are allow parallel access of all 9 disks.

RAID5 access time =16 ms that is accessing all 9 disk and having one read is ataking 16 msec time.

In RAID 3 we have Byte striping so even accessing 128 Bytes on disks we need (128/8=16 read cycle).

Therefore accessing 1 KB of data will take 16 reads in RAID 3 where 1 read is 16 ms

Total access time of RAID 3 for 1 KB of data=16*16=256 ms.

The detailed explanation clearly showing why RAID 3 is taking large time then RAID 5 for same disk configuration and now a days RAID 3 is obsolete.